# OmniDocBench: Benchmarking Diverse PDF Document Parsing with Comprehensive Annotations

### v1

@@ Abstract
Document content extraction is crucial in computer vision, especially for meeting the high-quality data needs of large language models (LLMs) and retrieval-augmented generation (RAG) technologies. However, current document parsing methods suffer from significant limitations in terms of diversity and comprehensive evaluation. To address these challenges, we introduce OmniDocBench, a novel multi-source benchmark designed to advance automated document content extraction. OmniDocBench includes a meticulously curated and annotated high-quality evaluation dataset comprising nine diverse document types, such as academic papers, textbooks, slides, among others. Our benchmark provides a flexible and comprehensive evaluation framework with 19 layout category labels and 14 attribute labels, enabling multi-level assessments across entire datasets, individual modules, or specific data types. Using OmniDocBench, we perform an exhaustive comparative analysis of existing modular pipelines and multimodal end-to-end methods, highlighting their limitations in handling document diversity and ensuring fair evaluation. OmniDocBench establishes a robust, diverse, and fair evaluation standard for the document content extraction field, offering crucial insights for future advancements and fostering the development of document parsing technologies. The codes and dataset is available in https://github.com/opendatalab/OmniDocBench.
@@ 1 Introduction
Figure 1: OmniDocBench Data Diversity. It contains 9 PDF page types, along with Layout Annotations and Recognition Annotations. Furthermore, there are 5 Page Attributes, 3 Text Attributes, and 6 Table Attributes.
Document parsing is a foundational task in computer vision, focused on accurately extracting content from documents [18, 36, 39, 41, 45]. High-quality document content extraction typically involves the integration of multiple algorithmic modules. Layout detection algorithms identify different content areas on a page, OCR technology converts images of text regions into text, while formula and table recognition models identify specific regions and transform them into corresponding source code. These modules and reading order algorithms form a comprehensive process of converting documents into machine-readable formats.
With large models increasingly requiring high-quality data, the importance of document content extraction has become more pronounced. Although vast amounts of data are available online for training, knowledge-rich document data is relatively scarce. Documents such as academic papers and technical reports contain rich structured information that can significantly enhance the knowledge depth of large models. Moreover, the development of retrieval-augmented generation (RAG) [21, 10] technology relies on extracting accurate information from documents to improve the quality and relevance of generated content. Consequently, research in document content extraction has intensified, leading to a series of pipeline-based high-quality document extraction algorithms [36] and the emergence of end-to-end multimodal large model solutions [5, 40, 27, 39, 3, 6, 42]. These methods have significantly improved document content parsing quality, providing robust support for the needs of large models and RAG technology.
In analyzing current module-based pipeline and multimodal end-to-end methods, we identified several limitations. For instance, methods like Marker and MinerU, which are mainstream pipeline methods, primarily evaluate individual modules on academic paper data, lacking document diversity and comprehensive evaluation results. Although MinerU considers the generalization of diverse data, it only demonstrates this through a single model and visualization results, lacking overall end-to-end evaluation. Multimodal large model methods [5, 40, 27, 39, 3], while easier to use than pipeline methods, lack performance validation on diverse documents, and some evaluation metrics are inadequate. Additionally, these methods often use data similar to their training set distribution for comparison, resulting in unfair evaluations. Overall, current document content extraction faces the following challenges:
  • Limited document types. Current evaluations mostly focus on a single type of academic paper, while real-world scenarios include textbooks, exam papers, financial reports, newspapers, magazines, and other document types.
  • Monotonous evaluation dimensions. Pipeline-based methods typically evaluate specific algorithmic modules, such as OCR, layout detection, or formula recognition, while the overall quality of parsing results requires comprehensive metrics.
  • Inadequate evaluation metrics. Multimodal large model approaches attempt to evaluate document parsing quality across multiple dimensions, such as dividing document content into text, formulas, tables, etc. However, these models commonly employ evaluation metrics such as BLEU scores or Edit distances, which fail to accurately and fairly assess parsing effectiveness when dealing with markup languages like LaTeX or HTML that allow diverse syntactic expressions.
Building a diverse, comprehensive, and accurate evaluation system poses significant challenges, requiring diverse and high-quality data annotation and reasonable evaluation metrics. While READOC extends the evaluation scope to include GitHub README files based on Nougat, there remains a substantial gap in real-world diversity, and the evaluation dimensions lack consideration of attributes. In contrast, this paper proposes a document content extraction benchmark, OmniDocBench, characterized by diverse types, detailed annotations, and reasonable evaluation (Figure 1). The specific contributions are as follows:
  • High-quality, diverse evaluation set: Through automated annotation, manual verification, and expert review, we construct a comprehensive, detailed, high-quality OmniDocBench evaluation set, encompassing nine types of diverse document pages, including papers, textbooks, exam questions, and research reports.
  • Flexible and comprehensive evaluation dimension support: The OmniDocBench validation set covers 19 layout category labels and 14 attribute labels. To facilitate user evaluation from an overall, single module, or different data types, we provide end-to-end evaluation, single algorithm module evaluation, and attribute-based evaluation, covering various evaluation needs.
  • Comprehensive evaluation of mainstream methods: Based on OmniDocBench, we conduct a comprehensive evaluation of current mainstream modular pipeline and end-to-end large model methods, providing a fairer assessment of existing methods and summarizing the shortcomings of current document parsing methods, thereby guiding further development in document parsing.
@@ 2 Related Work
Benchmark Document Categories Annotaion Type Single-Module Eval End-to-End Eval
BBox Text Table Order Formula OCR DLA TR MFR OCR ROD TR MFR
Single-Module Eval Benchmark
@@
PubTabNet [47],TableX [9], TableBank [23] 1, 1, 1
@@
Table 1: A comparison between OmniDocBench and existing DCE benchmarks. OCR: Optical Character Recognition; DLA: Document Layout Analysis; MFR: Math Formula Recognition; TR: Table Recognition; ROD: Reading Order Detection
2.1 Traditional Document Content Extraction
Document Content extraction remains a challenging task, and there is yet to emerge a unified benchmark tailored for real-world scenarios. Traditional algorithms typically employ multiple expert modules to handle different extraction subtasks, such as document layout detection [16, 46, 32, 11], optical character recognition (OCR) [33, 22, 28, 14, 37], formula recognition [44, 25, 4, 34], and table recognition [15, 17, 22].
While expert models of these subtasks are advancing rapidly, recent work such as Mineru [36] attempts to concatenate multiple expert modules into a pipeline and provides a high-precision open-source solution for document content extraction. READOC [26] also unifies heterogeneous evaluation methods from the perspective of Document Structure Extraction, breaking down texts, images, formulas, tables, and other dimensions for evaluation, thus offering a solution-oriented towards real-world scenarios for DSE tasks. However, due to the complexity of Document data sources and the intricacies of PDF document information, previous efforts still fall short in terms of data diversity, failing to cover the categories users encounter in practical applications. Similarly, there is an issue with the explainability of document parsing.
@@ 2.2 VLM-based Document Content Extraction
The emergence of Vision-Language Models (VLMs) [38, 6, 1, 12] has revolutionized the field of document content extraction. These models leverage multi-modality capability to achieve remarkable performance in document understanding tasks. Document extraction tools powered by VLMs excel at comprehending both visual layouts and textual content, effectively handling complex document structures while capturing rich contextual information. Representative works such as Nougat [5], Vary [40], Fox [27], and GOT [39], along with recent advances [13, 29], demonstrate significant progress in automated document parsing and comprehension. Despite these advances, the field lacks a standardized and unified benchmark for evaluating VLM-based document extraction task. This absence has hindered objective assessment of PDF document processing capabilities and impeded fair comparison across different approaches. To address this limitation, we present OmniDocBench, a comprehensive end-to-end benchmark designed specifically for evaluating VLM-based document parsing in real-world scenarios.
2.3 Benchmark for Document Content Extraction
An end-to-end benchmark for PDFs can intuitively reflect the effectiveness of PDF extraction tools, which is crucial for their iteration and selection. However, current benchmarks predominantly focus on module-level evaluations; we have listed related benchmarks in Table 1. Additionally, while there are existing end-to-end benchmarks, they lack detailed annotation rules and suffer from insufficient diversity, as well as unreasonable metrics for formula and table evaluations. For example, READOC [26] covers only two types of sources—arXiv and GitHub—and uses EDS [20] and TEDS [47] to compute metrics for formulas and tables, which may lead to inaccuracies CDM [35]. Therefore, there is a need for a more finely annotated, diverse, and reasonably evaluated end-to-end benchmark.
@@ 3 OmniDocBench Dataset
Constructing a diverse and comprehensive document parsing benchmark with precise annotations is a formidable challenge. As illustrated in Figure 2, we have designed a systematic and professional annotation framework for OmniDocBench, encompassing data acquisition, intelligent pre-annotation, and manual refinement. This ensures that OmniDocBench possesses the following key attributes:
@@
  • Comprehensive Annotation. We meticulously annotated all elements on the pages, including bounding boxes, specific content, and various potential attributes.
@@ 3.1 Data Acquisition
During the data acquisition phase, we sourced document pages from diverse origins and used clustering algorithms to initially select visually diverse pages, followed by manual annotation of page attributes to finalize the OmniDocBench pages. Specifically, we collected 200,000 initial PDF documents from Common Crawl, Google, Baidu search engines, and internal data. Subsequently, we extracted visual features from these document pages using ResNet-50 and performed clustering using Faiss  [^1] https://github.com/facebookresearch/faiss, sampling 6,000 visually diverse pages from 10 cluster centers. Finally, annotators provided page-level attribute annotations, including page type, layout type, and language type, and further balanced the selection to 981 samples for the final dataset. The OmniDocBench dataset includes pages with 9 types of pages, multiple layout categories, and various attribute annotations, covering a wide range of real-world scenarios.
@@ 3.2.1 Annotation Types
Layout Detection Annotations: Unlike typical layout detection tasks, OmniDocBench includes four comprehensive types of annotations: (1) Layout Bounding Box Annotations: Locating information for 19 types of regions such as titles, text paragraphs, tables, and images. (2) Layout Attribute Annotations: Detailed attribute annotations for detected boxes, including 3 text box attributes, 6 table attributes, and 2 formula attributes. (3) Reading Order Annotations: Annotating the reading sequence of detected boxes. (4) Affiliation Annotations: For images, tables, formulas, and code blocks, we annotate captions and titles to distinguish them from main text. Similarly, for cross-page paragraphs, we annotate affiliation relationships.
Content Recognition Annotations: Based on the format of the content area, we conduct the following three types of area annotations: (1) Text Annotations: Pure text annotations for titles, text paragraphs, and other plain text content. (2) Formula Annotations: LaTeX format annotations for inline formulas, display formulas, and subscripts. (3) Table Annotations: Providing both HTML and LaTeX annotations for table data.
@@ 3.2.2 Annotation Process
For these annotation tasks on diverse pages, we design a standardized process to ensure quality and efficiency, comprising intelligent pre-annotation, annotator correction, and expert quality inspection.
Intelligent Pre-Annotation. Manually annotating entire documents is time-consuming and costly. To enhance efficiency, we employ state-of-the-art detection and recognition models for pre-annotation of layout detection and content recognition. Specifically, we use fine-tuned LayoutLMv3 [16] for layout detection annotations and PaddleOCR [22], UniMERNet [34], and GPT-4o [2] for text, formula, and table annotations, respectively.
Annotator Correction. After layout detection phase, annotators refine the detection boxes and enhance annotations with reading order and affiliation details. Each character is verified to ensure accuracy in content recognition. For complex annotations of tables and formulas, requiring LaTeX and HTML formats, annotators use tools like Tables Generator  [^2] https://www.tablesgenerator.com/ and latexlive  [^3] https://www.latexlive.com/ for verification and correction.
Expert Quality Inspection. Despite thorough annotator corrections, the complexity of formulas and tables may result in residual issues. To address these, we use CDM’s rendering techniques to identify unrenderable elements. These are then reviewed and corrected by three researchers to ensure accuracy and fidelity in the final annotations.
@@ 3.3 Dataset Statistics
Annotation Diversity: OmniDocBench contains over 10,0000 annotations for page detection and recognition: (1) More than 20,000 block-level annotations across 15 categories, including over 9,000 text paragraphs, 989 image boxes, 428 table boxes, and so on. All document components except headers, footers, and page notes are labeled with reading order information, totaling over 16,000 annotations. (2) The dataset also includes more than 80,000 span-level annotations across four categories, with 4,000 inter-line formulas and footnote markers represented in LaTeX format, while the remaining annotations are in text format.
Annotation Attribute Diversity: (1) Text Attributes: All block-level annotations, except for tables and images, include text attribute tags. In addition to standard Chinese and English text, there are over 2,000 blocks with complex backgrounds and 146 with rotated text. (2) Table Attributes: Besides standard Chinese and English tables, there are 142 with complex backgrounds, 81 containing formulas, 150 with merged cells, and 7 vertical tables.

### v2

@@ Abstract
Document content extraction is a critical task in computer vision, underpinning the data needs of large language models (LLMs) and retrieval-augmented generation (RAG) systems. Despite recent progress, current document parsing methods have not been fairly and comprehensively evaluated due to the narrow coverage of document types and the simplified, unrealistic evaluation procedures in existing benchmarks. To address these gaps, we introduce OmniDocBench, a novel benchmark featuring high-quality annotations across nine document sources, including academic papers, textbooks, and more challenging cases such as handwritten notes and densely typeset newspapers. OmniDocBench supports flexible, multi-level evaluations—ranging from an end-to-end assessment to the task-specific and attribute-based analysis—using 19 layout categories and 15 attribute labels. We conduct a thorough evaluation of both pipeline-based methods and end-to-end vision-language models, revealing their strengths and weaknesses across different document types. OmniDocBench sets a new standard for the fair, diverse, and fine-grained evaluation in document parsing. Dataset and code are available at https://github.com/opendatalab/OmniDocBench.
@@ 1 Introduction
Figure 1: Results of End-to-End Text Recognition on OmniDocBench across 9 PDF page types.
Figure 2: Overview of OmniDocBench Data Diversity. The benchmark includes 9 diverse PDF document types. It supports rich annotation types, including layout annotations (e.g., title, table, figure) and recognition annotations (e.g., text spans, equations, tables). Each page is annotated with 6 page-level attributes (e.g., PDF type, layout type), along with fine-grained 3 text attributes (e.g., language) and 6 tables attributes (Items under “special issues” are treated as individual binary attributes (yes/no)), enabling detailed and robust evaluation.
As large language models [1, 39, 44, 28] increasingly rely on high-quality, knowledge-rich data, the importance of accurate document parsing has grown substantially. Document parsing, a core task in computer vision and document intelligence, aims to extract structured, machine-readable content from unstructured documents such as PDFs. This task is particularly critical for ingesting academic papers, technical reports, textbooks, and other rich textual sources into large language models, thereby enhancing their factual accuracy and knowledge grounding [19, 42, 45, 47, 52]. Moreover, with the emergence of retrieval-augmented generation (RAG) systems [22, 12], which retrieve and generate answers conditionally with external documents, the demand for precise document understanding has further intensified.
To address this challenging task, two main paradigms have emerged: 1) Pipeline-based approaches that decompose the task into layout analysis, OCR, formula/table recognition, and reading order estimation [42, 34]; and 2) End-to-end vision-language models (VLMs) that directly output structured representations (e.g., Markdown) [7, 46, 29, 45, 3, 8, 48]. Although both approaches have demonstrated promising results, conducting a broad comparison of their effectiveness remains challenging due to the absence of a comprehensive and unified evaluation benchmark.
As shown in Table 1, for pipeline-based document parsing systems, dedicated benchmarks [26, 54, 10] have been developed to target specific sub-tasks. For end-to-end evaluation, works like Nougat [7] and GOT-OCR [45] provide relatively small validation sets and assess predictions using page-level metrics such as Edit Distance [21].
However, these benchmarks present several key limitations: 1) Limited document diversity: Existing datasets primarily focus on academic papers, overlooking other real-world document types such as textbooks, exams, financial reports, and newspapers; 2) Inconsistent evaluation metrics: Current benchmarks rely heavily on generic text similarity metrics (e.g., Edit Distance [21] and BLEU [33]), which fail to fairly assess the accuracy of formulas and tables in LaTeX or HTML formats that allow for diverse syntactic expressions; and 3) Lack of fine-grained evaluation: Most evaluations report only an overall score, lacking insights into specific weaknesses, such as element-level score (e.g., text vs. formula) or per document-type performance (e.g., magazine or notes).
To address these limitations, we introduce OmniDocBench, a new benchmark designed to provide a rigorous and comprehensive evaluation for document parsing models across both pipeline-based and end-to-end paradigms. In summary, our benchmark introduces the following key contributions:
  • High-quality, diverse evaluation set: We include pages from 9 distinct document types, ranging from textbooks to newspapers, annotated using a combination of automated tools, manual verification, and expert review.
  • Flexible, multi-dimensional evaluation: We support comprehensive evaluation at three levels—end-to-end, task-specific, and attribute-based. End-to-end evaluation measures the overall quality of full-page parsing results. Task-specific evaluation allows users to assess individual components such as layout detection, OCR, table recognition, or formula parsing. Attribute-based evaluation provides fine-grained analysis across 9 document types, 6 page-level attributes and 9 bbox-level attributes.
  • Comprehensive benchmarking of state-of-the-art methods: We systematically evaluate a suite of representative document parsing systems, including both pipeline-based tools and VLMs, providing the most comprehensive comparison and identifying performance bottlenecks across document types and content structures.
@@ 2 Related Work
Benchmark Document Domain Annotaion Type Single-Task Eval End-to-End Eval
BBox Text Table Formula Attributes OCR DLA TR MFR OCR TR MFR ROD
Single-Task Eval Benchmark
@@
PubTabNet [54],TableX [11] 1, 1
TableBank [25] 1
@@
Table 1: A Comparison between OmniDocBench and existing benchmarks. BBox: Bounding boxes. Text: Text in Unicode. Table: Table in LaTeX/HTML/Markdown. Formula: Formula in LaTeX. Attributes: Page- and BBox-Level Attributes. OCR: Optical Character Recognition; DLA: Document Layout Analysis; TR: Table Recognition; MFR: Math Formula Recognition; ROD: Reading Order Detection
2.1 Pipeline-based Document Content Extraction
Pipeline-based methods treat the document content extraction task as a collection of single modules, such as document layout detection [17, 53, 36, 13], optical character recognition [38, 23, 30, 15, 43], formula recognition [51, 27, 6, 40], and table recognition [16, 18, 23]. In this sense, such methods can utilize different expert models to address each specific task. Marker [34] integrates open-source models to parse documents into structured formats such as Markdown, JSON, and HTML. To get higher accuracy, an optional LLM-enabled version can also be integrated to merge tables across pages, handle inline math, and so on. Similarly, MinerU [42] first utilizes a layout detection model to segment the document page into different regions, then applies task-specific models for corresponding regions. Finally, it outputs the complete content in Markdown format with a reading order algorithm. By leveraging lightweight models and parallelized operations, pipeline-based methods can achieve efficient parsing speeds.
@@ 2.2 VLM-based Document Content Extraction
Document understanding and optical character recognition (OCR) are crucial tasks for evaluating the perception capabilities of vision-language models (VLMs). By incorporating extensive OCR corpus into the pretraining stage, VLMs like GPT4o [2] and Qwen2-VL [3] have demonstrated comparable performance in document content extraction tasks. Unlike pipeline-based methods, VLMs perform document parsing in an end-to-end manner. Furthermore, without requiring specialized data fine-tuning, these models are able to deal with diverse and even unseen document types for their generalization capabilities.
To integrate the efficiency of lightweight models and the generalizability of VLMs, many works [7, 46, 29, 45, 14, 32] have focus on training specialized end-to-end expert models for document parsing. These VLM-driven models excel at comprehending both visual layouts and textual contents, balancing a trade-off between accuracy and efficiency.
2.3 Benchmarks for Document Content Extraction
Document content extraction requires the ability to understand document layouts and recognize various types of content. However, current benchmarks fall short of a comprehensive page-level evaluation, as they focus solely on evaluating the model’s performance on module-level recognition. PubLayNet [49] and concurrent benchmarks [26, 35, 9] specialize in evaluating a model’s ability to detect document page layouts. OCRBench [31] proposes five OCR-related tasks with a greater emphasis on evaluating the model’s visual understanding and reasoning capabilities. Only line-level assessments are provided for text recognition and handwritten mathematical expression recognition (HMER). Similarly, single-module benchmarks [20, 54, 26, 40] disentangle the task into different dimensions and focus narrowly on specific parts. Such paradigm overlooks the importance of structural and semantic information like the reading order and fails to evaluate the model’s overall ability when processing the full-page documents as a whole.
Page-level benchmarks have been proposed alongside some recent VLM-driven expert models [29, 7, 45]. However, the robustness of these benchmarks is compromised by limitations in data size, language, document type, and annotation. For example, Nougat [7] evaluates models using only printed English documents collected from arXiv while the page-level benchmark introduced by GOT-OCR [45] consists of only 90 pages of Chinese and English documents in total. Commonly-seen document types like handwritten notes, newspapers, and exam papers are further neglected. Lacking detailed annotations, the benchmarks can only conduct naive evaluation between the full-page results of Ground Truths and predictions without special handling for different output formats and specialized metrics for different content types. The evaluation of the model performance can be severely biased due to limited document domains, unaligned output format and mismatched metrics. Therefore, there is an urgent need for a more finely annotated, diverse, and reasonable page-level document content extraction benchmark.
@@ 3 OmniDocBench Dataset
Constructing a diverse and comprehensive document parsing benchmark with precise annotations is a significant challenge. As illustrated in Figure 3, we have designed a systematic and professional annotation framework for OmniDocBench, encompassing data acquisition, intelligent pre-annotation, and manual refinement. This ensures that OmniDocBench possesses the following key attributes:
@@
  • Comprehensive Annotation. We meticulously annotated all elements on the pages, including bounding boxes, specific contents, and various potential attributes.
@@ 3.1 Data Acquisition
During the data acquisition phase, we sourced document pages from diverse origins and used clustering algorithms to initially select visually diverse pages, followed by manual annotation of page attributes to finalize the OmniDocBench pages. Specifically, we collected over 200,000 initial PDF documents from Common Crawl, Google, Baidu search engines, and internal data. Subsequently, we extracted visual features from these document pages using ResNet-50 and performed clustering using Faiss  [^1] https://github.com/facebookresearch/faiss, sampling 6,000 visually diverse pages from 10 cluster centers. Finally, annotators provided page-level attribute annotations, including page type, layout type, and language type, and further balanced the selection to 981 samples for the final dataset. The OmniDocBench dataset includes pages from nine distinct types, multiple layout categories, and various attribute annotations, covering a wide range of real-world scenarios.
@@ 3.2.1 Annotation Types
Layout Detection Annotations: Unlike typical layout detection tasks, OmniDocBench includes four comprehensive types of annotations: (1) Layout Bounding Box Annotations: Positioanl information for 19 distinct region categories such as titles, text paragraphs, tables, and images. (2) Layout Attribute Annotations: Detailed attribute annotations for detected boxes, including 3 text box attribute categories, 6 table attribute categories, 9 bbox-level attribute labels in total. (3) Reading Order Annotations: Annotating the reading sequence of detected boxes. (4) Affiliation Annotations: For images, tables, formulas, and code blocks, we annotate captions and titles to distinguish them from main text. Similarly, for cross-page paragraphs, we annotate affiliation relationships.
Content Recognition Annotations: Based on the content type within each region, we conduct the following three types of annotations: (1) Text Annotations: Pure text annotations for titles, text paragraphs, and other plain text content. (2) Formula Annotations: LaTeX format annotations for inline formulas, display formulas, and subscripts. (3) Table Annotations: Providing both HTML and LaTeX annotations for table data.
@@ 3.2.2 Annotation Process
For these annotation tasks on diverse pages, we design a standardized process to ensure quality and efficiency, comprising intelligent automatic annotation, annotator correction, and expert quality inspection.
Automatic Annotation. Manually annotating entire documents is time-consuming and costly. To enhance efficiency, we employ state-of-the-art detection and recognition models for pre-annotation of layout detection and content recognition. Specifically, we use fine-tuned LayoutLMv3 [17] for layout detection annotations and PaddleOCR [23], UniMERNet [40], and GPT-4o [2] for text, formula, and table annotations, respectively.
Annotator Correction. After the layout detection phase, annotators refine the detection boxes and enhance annotations with reading order and affiliation details. Each character is verified to ensure accuracy in content recognition. For complex annotations of tables and formulas, requiring LaTeX and HTML formats, annotators use tools like Tables Generator  [^2] https://www.tablesgenerator.com/ and latexlive  [^3] https://www.latexlive.com/ for verification and correction.
Expert Quality Inspection. Despite thorough annotator corrections, the complexity of formulas and tables may result in residual issues. To address these, we use CDM’s rendering techniques [41] to identify unrenderable elements. These elements are then reviewed and corrected by three researchers to ensure accuracy in the final annotations.
@@ 3.3 Dataset Statistics
Annotation Diversity: OmniDocBench contains over 100,000 annotations for page detection and recognition: (1) More than 20,000 block-level annotations across 15 categories, including over 15,979 text paragraphs, 989 image boxes, 428 table boxes, and so on. All document components except headers, footers, and page notes are labeled with reading order information, totaling over 16,000 annotations. (2) The dataset also includes more than 70,000 span-level annotations across 4 categories, with 4,009 inline formulas and 357 footnote markers represented in LaTeX format, while the remaining annotations are in text format.
Annotation Attribute Diversity: (1) Text Attributes: All block-level annotations, except for tables and images, include text attribute tags. In addition to standard Chinese and English text, there are over 2,000 blocks with complex backgrounds and 493 with rotated text. (2) Table Attributes: In addition to standard Chinese and English tables, there are 142 tables with complex backgrounds, 81 containing formulas, 150 with merged cells, and 7 vertical tables.